First visit after lockdown.   The custodians and Rowan, attended to go through the sterilization of all apiary equipment that was likely to come into contact with bees from colony 3 and to burn and clean equipment from the Warre hive that was culled due to the possibility that a weak colony could easily catch EFB.  Rowan did a grand job with the fire ( day off for a firefighter!).
Temperature 25 degrees C  and Sunny.   No wind.    –  A lovely day.
Hive#1:   Inspected by Richard.  Brood Chamber:  14 x 12.
Queen found at last and Killed.(Regicide), due to high levels of chalk brood.  No Qc’s.   6 f BAS, stores and room ok, 1 super starting to fill,     Check next week and reduce Qc’s to 2. Temper 9.
Hive#2:    Inspected by Richard.   Brood chamber   14 x 12.
Queen seen,  2 qc’s taken down,   8 f BAS,  stores and room OK,  1 super Part filled. temper 9.  Bald brood still evident. signs of varroa damage.  big colony.
Hive#3:    Culled due to EFB.
Hive#4:   Inspected by Richard,  Brood chamber: National Deep.
QNS,  5 sealed Qc’s.  7 f BAS and day old eggs. Stores and room ok,   3 supers, 1 full, 1 new foundation, bottom raised up from brood chamber with brood on, as advised by the SBI, as mainly stores.    Frame with Qc’s on used to make up another Nuc, along with a shallow frame of stores. To be replaced ASAP.  Deep frame replaced with foundation to give more room in brood nest.   Other Qc’s taken down.   Temper 9.
Hive#5:   Inspected by Liz.   Brood Chamber: Deep and a Half.
QNS,   10 Qc’s  play cups, some charged,  taken down .  6f BAS, stores and room OK. 2 supers.  lots of happy bees.   Temper 8.
Hive#6:  Inspected   by Liz.   Brood chamber: Brood and a half.
QNS,   8 Qc’s,  some play cups, taken down. 8 f BAS, +day old eggs. stores mainly in shallow brood box.   1 super, part filled.  Temper 7.
Hive#7:  Inspected by Mel.     Brood Chamber:   Deep.
QNS, no Qc’s.  no brood,  stores and room OK,  2 Supers , 1 Full.   No evidence of eggs yet.               This was an artificial swarm ( 18/5 ) .  with Q left on foundation and fed syrup.  is she still there?
Hive#7A:  Inspected by Mel.    Brood chamber:  4 Shallows.
QNS, no Qc’s.  some brood still but no new eggs. stores and room OK.   no supers.  Temper 7. bees very agitated.  Should have been a queen cell somewhere !
Hive#8:   Inspected by Mel.   Brood Chamber : Deep.
QNS,  12 QC’s lots of brood but no eggs.  ( Q killed 29/5) stores and room OK.  1 super.
Hive#9:   Culled on the advice of SBI
Hive#10   Top Bar,   Not Inspected
Hive#15F  Inspected by Liz,    Brood Chamber: Deep.
QNS,  No Qc’s,   4.5 f  BAS,  stores and room ok.  No Supers.  temper 9.   some play cups. bees on frames 6 and 7,   It was hoped to use this colony in the observation hive for the coming two fete’s but due to the lock down we are unable.   We will now make up to a full colony.
Hive#15H     Not Inspected.
Hive#16A     NUC made up last week with Queen cell, not inspected, but bees still flying a good sign.